Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 456     Entry time: 23 Nov 2023, 22:15
Author: Jörg Hallmann 
Group: MID 
Subject: Shift summary 

progress with the Damnit - XPCS analysis test with various samples (SiO2, Ludox etc.)

tested some external samples P3HT, SOD1 but it looks like conventional XPCS is not working (signal too low - increasing transmission leads to radiation damage in both cases)

checked the diamond attenuators in the tunnel & optics hutch - all thicknesses indicated in the scenes seem to be correct

 

issue: AGIPD HV switched off by itself - ramping up again was possible

ELOG V3.1.4-7c3fd00